# A Survey on Cross-Domain Sequential Recommendation

### v2

@@ 4.1.3 Graph Neural Networks in CDSR
[…]
@@
Table 4: Experimental results (%) on two domains of Foods and Kitchen of Amazon. Notice that the results are copied from Cao et al. (2022); Xiao et al. (2023); Ye et al. (2023) for reference. We bold the best results and underline the second-best results.
@@ 4.2.1 Transfer Learning
Transfer learning is primarily employed to transfer knowledge learned from one task to another task. For instance, CDNST Zhuang et al. (2020) transfers the novelty-seeking trait learned from a source domain to a target domain. SATLR Liu and Zhu (2021) considers transferring knowledge by multiplying the independently learned representations from one domain with an orthogonal mapping matrix. To improve the ability of knowledge transfer, some works adopt dual transfer learning. DAT-MDI Chen et al. (2021) combines a dual transfer model with slot attention to self-adapte item embedding from different domains. DASL Li et al. (2021) applies a dual embedding component to unify the learning process of user representations and then proposes a dual attention component to incorporate user behaviors in multiple domains.
@@ 4.2.2 Contrastive Learning
Contrastive learning (CL) is also a widely applied technique that leverages the similarities and differences between samples to extract useful information. In the case of Tri-CDR Ma et al. (2023), it designs two contrastive learning tasks, i.e., the coarse-grained similarity modeling and the fine-grained distinction modeling. The coarse-grained similarity modeling closes three domains’ sequence representations of the same user, and the fine-grained distinction modeling assumes the distance between domain A and domain B should be larger than the distance between domain B and the hybrid domain. MGCL Xu et al. (2023c) views the local and global item representations of a user as the positive samples and the representations from different users as the negative samples. Additionally, some works aggregate the sequences and then combine those processed sequences with CL. DREAM Ye et al. (2023) proposes supervised contrastive learning to minimize the relevance among inter-sequences with different preferences.
@@ 4.2.4 Discussion
Incorporating auxiliary learning technologies aims to enhance the model’s ability to capture cross-domain information. Contrastive learning can make learned representations more discriminative and robust, but it is sensitive to the designed contrastive strategy. Transfer learning enables the sharing of information across domains, but the effectiveness of knowledge transfer is affected by the correlation between domains, i.e., there is vulnerability to negative transfer. There is also adversarial learning that can unify representations from different domains but the training process is more complex.
@@ 5 Datasets and Experimental Results
In this section, we summarize a list of commonly used datasets in the CDSR scenario, including their corresponding domains, data types, and scales shown in Table 3.
In order to more fully represent the performance of models with different granularity of information and different technologies in CDSR, we quote the results of representative models from Cao et al. (2022); Xiao et al. (2023); Ye et al. (2023) in Table 4. Notice that these three papers are consistent in their treatment of this dataset. We can see that with the increase in information and advancements in methods, the results gradually improve.
@@ 6 Future Directions
Although certain progress has been made in addressing the CDSR problem, there is still ample room for improvement. In this section, we provide several promising directions in CDSR for potential developments from different aspects.
@@ 6.0.2 Heterogeneous Information Fusion
Apart from the side information mentioned in Section 2.3, there is a large amount of relevant heterogeneous information in the real world. For example, users are likely to transfer from browsing short videos to an e-commerce platform to purchase items mentioned in the videos they watched. Therefore, it is worth investigating effective methods that combine heterogeneous information (e.g., image, video, etc.) and traditional ID-based information, to address the challenges in cross-domain recommendation. Moreover, it is also worth considering introducing heterogeneous sequential behaviors Luo et al. (2022) into cross-domain scenarios.
@@ 6.0.3 Deep Utilization of Non-overlapping Information
Indeed, the majority of current models rely on overlapping users to bridge different domains, but non-overlapping data also contain hidden features that are worth exploring and extracting Liu et al. (2023). So researchers can conduct deeper studies on non-aligned information.
@@ 6.0.4 Privacy Preservation
When it comes to sensitive user information, encryption and protection of data is crucial. Particularly within the realm of CDSR, there is a greater inclusion of user data. Therefore, designing effective federated learning methods to ensure privacy while minimizing the loss of valuable information in the cross-domain scenario is a meaningful but less studied area Yang et al. (2019); Lin et al. (2023b).
@@ 6.0.6 More Advanced Technologies
In Section 4, we analyze the technologies used in existing CDSR models from a micro view. However, it is undeniable that it still requires more advanced technologies to propel models toward achieving greater leaps in performance. For instance, exploring the application of large language models Wu et al. (2023) in the CDSR scenario is also a promising direction.
7 Conlusions
Cross-domain sequential recommendation (CDSR) extends traditional recommender systems by incorporating sequential and cross-domain information, thus addressing data sparsity. In this survey, we approach CDSR with a four-dimensional tensor and offer a comprehensive overview from macro and micro views. From a macro view, we summarize the existing models by abstracting multi-level fusion structures. From a micro view, we focus on analyzing and summarizing the basic technologies and auxiliary learning technologies employed by the models. Finally, we also include some public datasets and experimental results for CDSR, and provide some promising future research directions.

### v3

@@ 4.1.3 Graph Neural Networks in CDSR
Aware of the structural relationships of item-item and item-user in a sequence, graph neural networks are used to model such information. Some works construct a graph per session Chen et al. (2021) or per domain Cao et al. (2022); Xu et al. (2023c), or just construct a global graph Guo et al. (2021); Zhang et al. (2023b) based on the hybrid domain of all users. Apart from that, DCDIR Bi et al. (2020) and MIFN Ma et al. (2022) construct a knowledge graph to encompass more semantic information. Most works apply graph encoders (i.e., GNNs) and sequence encoders (i.e., RNNs or attention mechanisms) together as complementary parts in CDSR modeling. For instance, DA-GCN Guo et al. (2021) utilizes graph convolutional networks to learn latent user representations and user-specific item representations and then carries the weights from the item and user neighbors to the target item in each domain with an attention matrix. C2DSR Cao et al. (2022) combines the encoded representations from GNNs with the original sequences and feeds them into the attention encoder for further modeling. It supplementally learns complex information about nodes and edges to capture more comprehensive preferences of users. However, when GNNs are applied to a large-scale data, the huge computational and storage overhead is a major drawback in most cases.
@@
Table 4: Experimental results (%) on two domains of the Foods and Kitchen of Amazon. Notice that the results are copied from Cao et al. (2022); Xiao et al. (2023); Ye et al. (2023) for reference. We bold the best results and underline the second-best results.
@@ 4.2.1 Transfer Learning
Transfer learning is primarily employed to transfer knowledge learned from one task to another task. For instance, CDNST Zhuang et al. (2020) transfers the novelty-seeking trait learned from a source domain to a target domain. SATLR Liu and Zhu (2021) considers transferring knowledge by multiplying the independently learned representations from one domain with an orthogonal mapping matrix. To improve the ability of knowledge transfer, some works adopt dual transfer learning. DAT-MDI Chen et al. (2021) combines a dual transfer model with slot attention to self-adapt item embedding from different domains. DASL Li et al. (2021) applies a dual embedding component to unify the learning process of user representations and then proposes a dual attention component to incorporate user behaviors in multiple domains.
@@ 4.2.2 Contrastive Learning
Contrastive learning (CL) is also a widely applied technique that leverages the similarities and differences between samples to extract useful information. In the case of Tri-CDR Ma et al. (2023), it designs two contrastive learning tasks, i.e., coarse-grained similarity modeling and fine-grained distinction modeling. Coarse-grained similarity modeling closes three domains’ sequence representations of the same user, and fine-grained distinction modeling assumes the distance between domain A and domain B should be larger than the distance between domain B and the hybrid domain. MGCL Xu et al. (2023c) views the local and global item representations of a user as the positive samples and the representations from different users as the negative samples. Additionally, some works aggregate the sequences and then combine those processed sequences with CL. DREAM Ye et al. (2023) proposes supervised contrastive learning to minimize the relevance among inter-sequences with different preferences.
@@ 4.2.4 Discussion
Incorporating auxiliary learning technologies aims to enhance a model’s ability to capture cross-domain information. Contrastive learning can make learned representations more discriminative and robust, but it is sensitive to the designed contrastive strategy. Transfer learning enables the sharing of information across domains, but the effectiveness of knowledge transfer is affected by the correlation between domains, i.e., there is vulnerability to negative transfer. Moreover, adversarial learning can unify representations from different domains, but its training process is more complex.
@@ 5 Datasets and Experimental Results
In this section, we summarize a list of commonly used datasets in CDSR, including their corresponding domains, data types, and scales shown in Table 3.
In order to more fully represent the performance of models with different granularities of information and different technologies in CDSR, we quote the results of some representative models from Cao et al. (2022); Xiao et al. (2023); Ye et al. (2023) in Table 4. Notice that these three papers are consistent in their treatment of this dataset. We can see that with the increase in information and advancements in methods, the results gradually improve.
@@ 6 Future Directions
Although certain progress has been made in addressing the CDSR problem, there is still ample room for improvement. In this section, we provide several promising directions in CDSR for potential development from different aspects.
@@ 6.0.2 Heterogeneous Information Fusion
Apart from the side information mentioned in Section 2.3, there is a large amount of relevant heterogeneous information in real-world applications. For example, users are likely to transfer from browsing short videos to purchasing items mentioned in the videos. Therefore, it is worth investigating effective methods that combine heterogeneous information (e.g., image, video, etc.) and traditional ID-based information, to address the challenges in cross-domain recommendation. Moreover, it is also worth considering introducing heterogeneous sequential behaviors Luo et al. (2022) into cross-domain scenarios.
@@ 6.0.3 Deep Utilization of Non-overlapping Information
Indeed, the majority of current models rely on overlapping users to bridge different domains, but non-overlapping data also contain rich semantic information that are worth exploring and extracting Liu et al. (2023). So researchers can conduct deeper studies on non-aligned information.
@@ 6.0.4 Privacy Preservation
When it comes to sensitive user information, data encryption and protection is crucial. Particularly within the realm of CDSR, there is a greater inclusion of user data. Therefore, designing effective federated learning methods to ensure privacy while minimizing the loss of valuable information in the cross-domain scenario is a meaningful but less studied area Yang et al. (2019); Lin et al. (2023b).
@@ 6.0.6 More Advanced Technologies
In Section 4, we analyze the technologies used in CDSR from a micro view. However, it is undeniable that more advanced technologies are still required to propel models toward achieving greater leaps in performance. For instance, exploring the application of large language models (LLMs) Wu et al. (2023) in CDSR is also a promising direction.
7 Conclusions
Cross-domain sequential recommendation (CDSR) extends traditional recommender systems by incorporating sequential and cross-domain information to address the data sparsity issue. In this survey, we approach CDSR with a four-dimensional tensor and offer a comprehensive overview from macro and micro views. From a macro view, we summarize the existing models by abstracting multi-level fusion structures. From a micro view, we analyze and summarize the employed basic and auxiliary learning technologies. Finally, we include some public datasets and experimental results for CDSR, and provide some promising future research directions.